The price for a good has reduced from 820 rub. to 790 rub. The demand has increased from 2900 units to 3350 units.. Calculate the price elasticity of demand? 


1
Expert's answer
2021-09-22T09:48:05-0400

"Price Elasticity of Demand = \frac{percentage change in quantity demanded}{percentage change in price}""Change in quantity demanded = 3350-2900 = 450"

%change in quantity demanded = "\frac{450}{2900 }×100 = 15.52"

"Change in price = 790-820 = -30"

Percentage change in price = "\frac{-30}{820}×100 =- 3.66"


"Price elasticity of demand = \frac{15.52}{-3.66}"

"= -4.24"